A jar of quarters and nickels contain 1.25. There are 13 coins in all. How many of each are there?
Answer by mananth(16949)

A jar of quarters and nickels contain 1.25. There are 13 coins in all. How many of each are there?
let the number of nickel coins be x
the quarters will be 13-x
value of nickels = 5x
Value of quarters = 25(13-x)
The total value 125 cents
5x+25(13-x)=125
5x+325-25x=125
-20x=-200
x=-200/-20
x=10 the number of nickel coins
13-10= 3 is the number of quarter coins
Ananth